OpenStudy (anonymous):

Write the quadratic function in vertex form. f(x) = x2 - 3x - 10

OpenStudy (callisto):

f(x) = x^2 - 3x - 10 = (x^2 - 3x + (3/2)^2 ) - (3/2)^2 -10 = (x-3/2)^2 - 49/4